MILES TRAVELED: GALLONS OF GAS USED:
22 1
44 2
88 4
154 7
330 15
QUESTION:
How far can Mr. Blake travel using 40 gal of gas?

Answer by Theo(13342)

he gets 22 miles to the gallon.
40 gallons * 22 miles per gallon = 40 * 22 = 880 miles.

the formula to use is:

number of miles driven = number of gallons used * number of miles per gallon.

you are getting 22 miles per gallon.

40 * 22 = 880.